OpenStudy (anonymous):

Of this quadratic equation, does a=-2, b=3, and c=-4?

OpenStudy (anonymous):

\[-2y ^{2}+3y=-4\]

OpenStudy (anonymous):

c = 4

OpenStudy (anonymous):

thanks!makes that become positive?

OpenStudy (anonymous):

-2y^2 + 3y +4 =0

OpenStudy (anonymous):

and would you mind if I try and solve it and post it on here so you can let me know if the answer I get is correct? :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[-3 + or - \sqrt{41} \over -4\] is this my answer or is there more I can do with it? Also, would there be 2 real answers?

OpenStudy (anonymous):

or 2 non real solutions since there are negative numbers..

OpenStudy (anonymous):

D = b^2-4ac = 9-(4*-2*4) =9+32 = 41 >0 means 2 real answer u are correct

OpenStudy (anonymous):

okk thanks. I'm kind of confused though. Would the 2 answers be what I put? If so, wouldn't they not be real since there are negative numbers? Sorry. I guess to clarify could you post what the final 2 answers are? :)

OpenStudy (anonymous):

(-3 - sqrt(41))/-4 and (-3 + sqrt(41))/-4 not real or imaginary is if sqrt(-41) - sqrt(41) is real because negative sign is out of sqrt

OpenStudy (anonymous):

oohh ok. thanks!

OpenStudy (anonymous):

no problem